OpenStudy (anonymous):

Someone help me please, I have been stuck on this question for over an hour and asked this 41 minutes ago. I have to finish this question and 9 others in 1 hour which is when they are due. The gross income of a business over time can be represented as a set of ordered pairs in which the first element, x, is the number of the year and the second element, y, is the gross income. The gross income is given for years 1 and 2. A constant rate of change is expected for the next several years. Year 1 (1997) Gross income (in 1000's) = $21 Year 2 (1998) Gross income (in 1000's) = $9 I have done the first two questions it has asked which are below: The ordered pair that represents the year and gross income for 1997 is (1,21). The ordered pair that represents the year and gross income for 1998 is (2,9). Here is the question I am stuck on. The linear equation, y=mx+b, relating gross income and years is y=__x+___.

hero (hero):

utilizing y = mx + b: 21 = 1m + b 9 = 2m + b Isolating b in each equation yields: b = 21 - 1m b = 9 - 2m setting b = b yields: 21 - 1m = 9 - 2m Placing like terms on the same side yields: 2m - 1m = 9 - 21 Solving for m gives: m = -11 Solving for b gives: b = 21 -1(-11) b = 21 + 11 b = 32 Therefore y = -11x + 32